OpenStudy (anonymous):

subtract rad44-2rad11

OpenStudy (anonymous):

\[\sqrt{44}-2\sqrt{11}=\sqrt{2^2*11}-2\sqrt{11}\]\[2\sqrt{11}-2\sqrt{11}=0\]